9. A 6.0 kg object (at rest) is released from a height of 80 m. If air resistance is negligible and g = 9.8 N/kg, the kinetic energy of the mass when it is 20 m above the ground is:

(a) 4.7 × 103 J (b) 3.5 × 103 J

(c) 2.4 × 103 J (d) 1.2 × 103 J



1
Expert's answer
2022-04-17T17:08:29-0400

According to the energy conservation law, the total energy at 20 m level its equal to the initial potential energy:


"mgh_0 = mgh +K"

where "h_0 = 80m" is the initial height, "h = 20m" is the current height, and "K" is the kinetic energy at 20 m level. Expressing "K", obtain:


"K = mg(h_0-h) \\
K= 6.0kg \cdot 9.8 N/kg\cdot (80m - 20m) \approx 3.5\times 10^ 3J"

Answer. (b) 3.5 × 103 J.